OpenStudy (anonymous):

The value of 123−−−√ lies between which two consecutive integers?

OpenStudy (phi):

? can you use the equation editor, or explain the question better?

OpenStudy (anonymous):

it says

OpenStudy (phi):

I would write down some integers and their squares. until I find the pair that works. for example, I know 10 squared is 10*10 = 100 which is smaller than 123 if 11*11 is bigger than 123 then we found the answer. otherwise we keep going...

myininaya (myininaya):

is 11*11=121 or do i actually need to multiply that

OpenStudy (phi):

I assume "modern" students can't do 11*11 in their head (but I could be wrong)

myininaya (myininaya):

oh oops you said if sorry

OpenStudy (anonymous):

log 16 256=? what about this one?

OpenStudy (phi):

You should know \[ \sqrt{100} = \sqrt{10 \cdot 10} = 10 \] right ? so it makes sense that \(\sqrt{123} \) is bigger than 10 but we should test 11 and 12 also

OpenStudy (phi):

Did you answer the sqrt(123) question ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

11 and 12. i am on a knew question now

OpenStudy (phi):

ok. But please make it a new post.

OpenStudy (anonymous):

i posted on here already

OpenStudy (anonymous):

log 16 256=?

OpenStudy (mathstudent55):

You mean \(\large \log_{16} 256\) If so, you are looking for the exponent you need to raise 16 to, to get 256. Hint: Write 256 as a power of 16.